OpenStudy (anonymous):

simplify (3y+7)-(-4x+2y)=

OpenStudy (espex):

Stat by distributing your negative and then combine like terms.

OpenStudy (anonymous):

Okay here are the steps... (3y+7)-(-4x+2y)= First you distribute the negative. 3y+7+4x-2y= You then combine like terms 7x+7=2y Isolate the y y=7/2x+7/2 Then you do a little math to get rid of the 2 from the y and there's your answer! :) Let me know if you need any more help. :)